When Tim was 14 years old his father was 42 years old, which was three times Tim's age. Now he is twice Tim's age. How old is Tim?
Answer by fractalier(6550)

Call the number of years in the future that this happens, x.
Then we can write
42 + x = 2(14 + x)
42 + x = 28 + 2x
14 + x = 2x
x = 14 years in the future
Thus, Tim is 28 and dad is 56.